# Modified scattering for the three dimensional Maxwell-Dirac system

### v4

@@ 5.3. An improved bound for the wave equation inside the cone
However we have to be careful that we have (5.47) only inside the cone. So this is useful only in the region where $S\gg\lambda^{-1}$, while outside we have to rely on (5.48). Precisely, we consider a cutoff function $\chi$ which selects the region $t-r>\lambda^{-1}$, so that $\nabla\chi$ has size $\lambda$ and support in the region $t-r\approx\lambda$. Then applying an inverse derivative at frequency $\lambda$ we write
@@
$=$ $+\lambda^{-1}K_{\lambda}\ast\chi\nabla\mathscr{T}A_{\lambda}+\lambda^{-1}K_{\lambda}\ast(\nabla\chi)\mathscr{T}A_{\lambda}+\lambda^{-1}\nabla K_{\lambda}\ast(1-\chi)\mathscr{T}A_{\lambda}.$
@@
which decays rapidly away from $t-r\approx\lambda$. We have proved that
@@ 6.2. Wave packets for the Dirac equation
$=:$ $\ e^{i\phi_{\pm}}(\chi(y)\sigma^{\pm}(x/t)+\sigma_{1}^{\pm}(x/t)),$
where $\sigma_{1}^{\pm}$ has similar regularity and localization as the first term, but has size $(t^{2}-x^{2})^{\frac{1}{4}}$ smaller. Applying the linear Dirac operator to $\psi_{v}^{\pm}$ we cancel the expression $P_{x/t}^{\pm}(R_{2}^{\pm}+R_{4}^{\pm})$, obtaining instead additional error terms. These are all $(t^{2}-x^{2})^{-\frac{3}{4}}$ better than $\psi_{v}^{\pm}$ with a single exception, namely the counterpart of $R_{4}^{\pm}$, which we denote by $R_{5}^{\pm}$, obtaining
@@ 6.3. Asymptotic profiles via wave packet testing
where $\psi_{v}^{\pm}$ are as in (6.12) and $\sigma^{\pm}_{1}(v)$ and $\sigma^{\pm}_{2}(v)$ form an orthonormal base with respect to $\langle\cdot,\cdot\rangle_{H}$ of $V^{\pm}_{v}$.
In the following, we will drop the subscript on $\sigma_{j}^{\pm}$.
@@
Splitting into the $\pm$ components and testing against $\sigma^{\pm}$, it suffices to show that
$\langle\psi(t,vt),\sigma^{\pm}(v)\rangle_{H}=(t^{2}-(vt)^{2})^{\frac{3}{4}}e^{i\phi_{\pm}}\int_{H}\langle\psi,\psi_{v}^{\pm}\rangle_{H}\,dV_{H}+O(\varepsilon(t^{2}-r^{2})^{-\frac{7}{8}}t^{C\varepsilon}).$
@@
$\langle\psi(t,vt),\sigma^{\pm}(v)\rangle_{H}=(t^{2}-(vt)^{2})^{\frac{3}{4}}\int_{H}\langle\psi,\chi(y)\sigma^{\pm}+\sigma_{1}^{\pm}\rangle_{H}\,dV_{H}+O(\varepsilon(t^{2}-(vt)^{2})^{-\frac{7}{8}}t^{c\varepsilon}).$
@@
The uniform bound (6.22) is used to estimate directly the contribution of $\sigma_{1}^{\pm}$, which has size $(t^{2}-x^{2})^{-\frac{1}{4}}$ times $\sigma_{v}^{\pm}$ and support similar to $\chi$, which has size $(t^{2}-x^{2})^{-\frac{1}{4}}$. Then we can estimate
$Err^{1}$= $\ \left|(t^{2}-(vt)^{2})^{\frac{3}{4}}\int_{H}\chi(y)\langle\psi,\sigma_{1}^{\pm}\rangle_{H}\,dV_{H}\right|$
@@
this can be essentially thought of as the difference between $f$ and its local average on the $(t^{2}-x^{2})^{-\frac{1}{4}}$ scale, up to an error that can be included in $\sigma_{1}^{\pm}$.
@@
(6.26) $i\partial_{t}\rho^{\pm}(t,v)=v^{\alpha}A_{\alpha}\rho^{\pm}(t,v)+e^{\pm 2i\phi}A_{\alpha}\sqrt{1-v^{2}}P^{\pm}_{v}\gamma^{\alpha}\rho^{\mp}(t,v)+O(\varepsilon t^{-\frac{3}{2}+c\varepsilon}(1-v^{2})^{-\frac{1}{4}}).$
@@
$\langle(\rho^{\pm}(t_{2},v)-\rho^{\pm}(t_{1},v)),\sigma_{v}^{\pm}\rangle_{H}=\int_{H_{[t_{1},t_{2}]}}iA_{\alpha}\langle\gamma^{0}\gamma^{\alpha}\psi,\psi_{v}^{\pm}\rangle-i\langle\gamma^{0}\psi,f_{v}^{\pm}\rangle\,dxdt.$
@@
$t_{1}\frac{d}{dt_{1}}\langle\rho^{\pm}(t_{1},v),\sigma_{v}^{\pm}\rangle_{H}=$ $\ \int_{H_{t_{1}}}\frac{t^{2}-x^{2}}{{\sqrt{t^{2}+x^{2}}}}\left(iA_{\alpha}\langle\gamma^{0}\gamma^{\alpha}\psi,\psi_{v}^{\pm}\rangle-i\langle\gamma^{0}\psi,f_{v}^{\pm}\rangle\right)\,{d\sigma}$
@@ 6.4. Global bounds for $\psi$
$C^{in}=\{t>t_{0}(v):=(1-v^{2})^{-10}\}.$
@@ 8.2. The radiation profile of AA
$\|(1-v^{2})^{-\frac{1}{2}}a^{\mu}_{\infty}\|_{C^{2,\frac{1}{4}}(H)}\lesssim\varepsilon^{2},$
@@
which has a known, explicit expression
$K(v,v^{\prime})=K(d)=\frac{c_{3}}{\sinh(d/2)\cosh^{2}(d/2)},\qquad d=d_{H}(v,v^{\prime}),$
which is a special case of the hypergeometric function.

### v5

@@ 5.3. An improved bound for the wave equation inside the cone
However we have to be careful that we have (5.48) only inside the cone. So this is useful only in the region where $S\gg\lambda^{-1}$, while outside we have to rely on (5.49). Precisely, we consider a cutoff function $\chi$ which selects the region $t-r>\lambda^{-1}$, so that $\nabla\chi$ has size $\lambda$ and support in the region $t-r\approx\lambda^{-1}$. Then applying an inverse derivative at frequency $\lambda$ we write
@@
$=$ $\lambda^{-1}K_{\lambda}\ast\chi\nabla\mathscr{T}A_{\lambda}+\lambda^{-1}K_{\lambda}\ast(\nabla\chi)\mathscr{T}A_{\lambda}+\lambda^{-1}\nabla K_{\lambda}\ast(1-\chi)\mathscr{T}A_{\lambda}.$
@@
which decays rapidly away from $t-r\approx\lambda^{-1}$. We have proved that
@@ 6.2. Wave packets for the Dirac equation
$=:$ $\ e^{i\phi_{\pm}}(\chi(y)\sigma^{\pm}(x/t)+\underline{\sigma}^{\pm}(x/t)),$
where $\underline{\sigma}^{\pm}$ has similar regularity and localization as the first term, but has size $(t^{2}-x^{2})^{\frac{1}{4}}$ smaller. Applying the linear Dirac operator to $\psi_{v}^{\pm}$ we cancel the expression $P_{x/t}^{\pm}(R_{2}^{\pm}+R_{4}^{\pm})$, obtaining instead additional error terms. These are all $(t^{2}-x^{2})^{-\frac{3}{4}}$ better than $\psi_{v}^{\pm}$ with a single exception, namely the counterpart of $R_{4}^{\pm}$, which we denote by $R_{5}^{\pm}$, obtaining
@@ 6.3. Asymptotic profiles via wave packet testing
where $\psi_{v}^{\pm}$ are as in (6.12) and $\sigma^{\pm}_{1}(v),\sigma^{\pm}_{2}(v)$ are chosen to form an orthonormal base with respect to $\langle\cdot,\cdot\rangle_{H}$ of $V^{\pm}_{v}$ and so that the last bullet point below (6.8) is satisfied.
In the following, we will sometimes drop the subscript on $\sigma_{j}^{\pm}$.
@@
Splitting into the $\pm$ components and testing against $\sigma_{j}^{\pm}$, it suffices to show that
$\langle\psi(t,vt),\sigma_{j}^{\pm}(v)\rangle_{H}=(t^{2}-(vt)^{2})^{\frac{3}{4}}e^{i\phi_{\pm}}\int_{H}\langle\psi,\psi_{v}^{\pm}\rangle_{H}\,dV_{H}+O(\varepsilon(t^{2}-r^{2})^{-\frac{7}{8}}t^{C\varepsilon}).$
@@
$\langle\psi(t,vt),\sigma_{j}^{\pm}(v)\rangle_{H}=(t^{2}-(vt)^{2})^{\frac{3}{4}}\int_{H}\langle\psi,\chi(y)\sigma^{\pm}+\underline{\sigma}^{\pm}\rangle_{H}\,dV_{H}+O(\varepsilon(t^{2}-(vt)^{2})^{-\frac{7}{8}}t^{c\varepsilon}).$
@@
The uniform bound (6.22) is used to estimate directly the contribution of $\underline{\sigma}^{\pm}$, which has size $(t^{2}-x^{2})^{-\frac{1}{4}}$ times $\sigma^{\pm}(v)$ and support similar to $\chi$, which has size $(t^{2}-x^{2})^{-\frac{1}{4}}$. Then we can estimate
$Err^{1}$= $\ \left|(t^{2}-(vt)^{2})^{\frac{3}{4}}\int_{H}\chi(y)\langle\psi,\underline{\sigma}^{\pm}\rangle_{H}\,dV_{H}\right|$
@@
this can be essentially thought of as the difference between $f$ and its local average on the $(t^{2}-x^{2})^{-\frac{1}{4}}$ scale, up to an error that can be included in $\underline{\sigma}^{\pm}$.
@@
(6.26) $i\partial_{t}\rho^{\pm}(t,v)=v^{\alpha}A_{\alpha}\rho^{\pm}(t,v)+e^{2i\phi_{\pm}}A_{\alpha}\sqrt{1-v^{2}}P^{\pm}_{v}\gamma^{\alpha}\rho^{\mp}(t,v)+O(\varepsilon t^{-\frac{3}{2}+c\varepsilon}(1-v^{2})^{-\frac{1}{4}}).$
@@
$\langle(\rho^{\pm}(t_{2},v)-\rho^{\pm}(t_{1},v)),\sigma^{\pm}(v)\rangle_{H}=\int_{H_{[t_{1},t_{2}]}}iA_{\alpha}\langle\gamma^{0}\gamma^{\alpha}\psi,\psi_{v}^{\pm}\rangle-i\langle\gamma^{0}\psi,f_{v}^{\pm}\rangle\,dxdt.$
@@
$t_{1}\frac{d}{dt_{1}}\langle\rho^{\pm}(t_{1},v),\sigma^{\pm}(v)\rangle_{H}=$ $\ \int_{H_{t_{1}}}\frac{t^{2}-x^{2}}{{\sqrt{t^{2}+x^{2}}}}\left(iA_{\alpha}\langle\gamma^{0}\gamma^{\alpha}\psi,\psi_{v}^{\pm}\rangle-i\langle\gamma^{0}\psi,f_{v}^{\pm}\rangle\right)\,{d\sigma}$
@@ 6.4. Global bounds for $\psi$
$C^{in}=\{(t,x)\in(0,\infty)\times\mathbb{R}^{3}\colon t>t_{0}(v):=(1-v^{2})^{-10},\text{ where }v=x/t\}.$
@@ 8.2. The radiation profile of AA
(8.20) $\|(1-v^{2})^{-\frac{1}{2}}a^{\mu}_{\infty}\|_{C^{2,\frac{1}{4}}(H)}\lesssim\varepsilon^{2},$
@@
which has a known explicit expression, cf. [24],
$K(v,v^{\prime})=K(d)=\frac{1}{4\pi\sinh d}$
which is a special case of a hypergeometric function. Since in the Klein-Beltrami model we have
$d(0,v)=\frac{1}{2}\log\frac{1+|v|}{1-|v|},$
it follows that
$K(0,v)=\frac{1}{4\pi}\frac{\sqrt{1-|v|^{2}}}{|v|}.$
The behavior near the unit sphere can be used to prove the expected limit behavior of $(1-|v|^{2})^{-\frac{1}{2}}a^{\mu}_{\infty}(v)$ as $|v|\to 1$ (we omit the details).